### v2

@@ 5.4 LLMs/VLMs-based Methods
To delve deeper into the context of LLMs and VLMs tailored for AVs, we now focus on recent methods explicitly developed for key AV-related tasks such as perception, reasoning, captioning, trajectory prediction, and scene understanding. Tables 23 and 24 present a comprehensive overview of the models, highlighting their core architectures, input modalities, evaluation metrics, and specific innovations tailored for AV systems.
Table 23: Comparative analysis of LLM-based models for autonomous driving perception, reasoning, captioning, and scene understanding, highlighting their architectures, modalities, evaluation metrics, and key innovations.
Ref Year Name Dataset Architecture Modalities Metrics Task
[yang2025lidar] 2025 LiDAR-LLM NuScenes-QA Encoder Text, LiDAR Accuracy Reasoning
Innovation: Reframes 3D LiDAR scene understanding as a language modeling task using a position-aware transformer.
[tang2025bev] 2025 BEV-TSR NuScenes Encoder Text, Vision Accuracy Reasoning
Innovation: Proposes BEV-TSR for text-to-BEV scene retrieval using LLMs and shared cross-modal embeddings.
[peng2025lc] 2025 LC-LLM HighD Encoder Text F1, Accuracy Reasoning
Innovation: Introduces an LLM-based model for lane change prediction with interpretable chain-of-thought reasoning.
[lan2024traj] 2025 Traj-LLM NuScenes GPT-2, LoRA Text minADE, MR Reasoning
Innovation: LLM-based trajectory predictor using lane-aware probabilistic learning and no prompt engineering.
[xu2024drivegpt4] 2024 DriveGPT4 BDD-X Multimodal Text, Video BLEU4 Reasoning
Innovation: An interpretable end-to-end AV system using LLMs with multi-frame video and textual reasoning.
[liu2024improved] 2024 LLaVA1.5 VQA-v2 Multimodal Text, Video Accuracy, BLEU Reasoning
Innovation: Demonstrates powerful LLM design with simple LLaVA modifications, achieving high performance.
[chen2024driving] 2024 GPT-3.5 Generated Multimodal Text Accuracy Captioning
Innovation: Proposes an object-level LLM that fuses numeric vectors with language for context understanding.
[hu2023gaia] 2023 GAIA-1 Generated Encoder-Decoder Text, Video Cross-entropy Understanding
Innovation: Introduces a generative world model predicting future driving scenarios using video, text, and actions.
[elhafsi2023semantic] 2023 GPT-3 Generated Multimodal Text, Vision CLIPort Reasoning
Innovation: Introduces an LLM-based monitor to detect semantic anomalies in AVs through human-like reasoning.
[wang2023drivemlm] 2023 DriveMLM CARLA Town05 Multimodal LiDAR, Vision F1, BLEU4 Reasoning
Innovation: Bridges LLM decisions with vehicle control for closed-loop autonomous driving using multimodal inputs.
[jin2023adapt] 2023 ADAPT Generated Multimodal Text, Video BLEU-1 Reasoning
Innovation: Introduces a real-time driving captioner that explains vehicle actions and decisions in natural language.
[mao2023gpt] 2023 GPT-DRIVER NuScenes Multimodal LiDAR, Vision L2 Reasoning
Innovation: Reformulates motion planning as a language modeling task using GPT-3.5 to generate precise trajectories.
In [yang2025lidar], the authors use LLMs to interpret sparse outdoor 3D LiDAR data. The proposed LiDAR-LLM reformulates 3D scene understanding tasks into a language modeling paradigm, such as captioning, grounding, and question answering. The authors design a three-stage training pipeline involving cross-modal alignment, object-centric perception, and high-level instruction tuning. The Position-Aware Transformer is central to the model, which integrates bird’s-eye view positional embeddings to align 3D LiDAR features with LLM token space. They also construct large-scale synthetic datasets (420K captioning and 280K grounding pairs) using nuScenes and show that LiDAR-LLM outperforms 2D vision-language baselines in multiple 3D tasks, achieving strong results on both their synthetic benchmarks and real-world datasets like NuScenes-QA. This work marks a significant step toward general-purpose, instruction-following LLMs for 3D scene understanding in AVs. The method proposed in the [lan2024traj] proposes a trajectory prediction in autonomous driving using pre-trained Large Language Models without explicit prompt engineering. The model captures high-level semantic interactions and scene context by encoding spatial-temporal features of agents and traffic scenes into LLM formats. It incorporates a novel lane-aware probabilistic learning module, powered by the Mamba architecture, to simulate human-like attention to lane structures. Moreover, a multi-modal Laplace decoder generates diverse and scene-compliant future trajectories. The results show that this method outperforms state-of-the-art baselines across key metrics, including in few-shot settings. It highlights its robustness, generalizability, and the untapped potential of LLMs for structured prediction in complex driving environments. [sima2024drivelm] presents a method that integrates VLMs into end-to-end autonomous driving by modeling human-like reasoning through Graph Visual Question Answering. This method captures multistep, multiobject dependencies by structuring perception, prediction, and planning tasks as interconnected Question Answering pairs in a graph. Moreover, the authors propose two datasets, DriveLM-nuScenes and DriveLM-CARLA, and a baseline model, DriveLM-Agent, which combines graph prompting with trajectory tokenization to repurpose general VLMs for driving tasks. Results demonstrate that this method performs well with driving-specific models and excels in zero-shot generalization, especially with unseen sensor configurations. This work highlights Graph VQA as a promising approach for enabling explainable, generalizable, and interactive AV system.
Table 24: Comparative analysis of VLM-based models for autonomous driving perception, reasoning, captioning, and scene understanding, highlighting their architectures, modalities, evaluation metrics, and key innovations.
Ref Year Name Backbone Dataset Modalities Metrics Task
[zhou2025dynrsl] 2025 DynRsl-VLM ViT+ResNet+Transformer NuInstruct Text, Image Accuracy, BLEU Reasoning
Innovation: Using dynamic resolution and efficient image-text alignment to preserve crucial visual details for driving VQA.
[zhou2025hierarchical] 2025 AV-VLM GPT-4V Generated Text, Vision, LiDAR Accuracy, Latency Reasoning
Innovation: Introduces a real-world test platform for evaluating VLM-driven autonomous systems under controlled, diverse driving scenarios.
[taparia2025vlc] 2025 VLC Fusion BEVFusion + CLIP Waymo, ATR Text, Vision, LiDAR IR, mAP Detection
Innovation: Using VLMs to adaptively weight sensor modalities based on environmental context cues for robust multimodal perception tasks.
[shriram2025towards] 2025 COOOL LLaMA-2 + CLIP COOOLER Video, Text BESM, SAM Captioning
Innovation: Introduces a multimodal VLM-LLM pipeline with CLIP for zero-shot hazard detection and explanation in traffic scenes.
[wu2025language] 2025 NuPrompt Transformer NuPrompt Text, Vision AMOTA, ADE Prediction
Innovation: An object-centric language prompt benchmark for 3D multi-view driving scenes with trajectory prediction.
[he2024integrating] 2024 Llama-Adapter Multimodal-7B Generated Text, Vision Accuracy, BLEU Reasoning
Innovation: Extends VLMs with YOLOS-based object detection and camera ID-separators for improved multi-view scene understanding.
[huang2024vlm] 2024 VLM-RL LLaMA+CLIP Generated Vision, Text AS, RC Reward Design
Innovation: Introduces VLM-RL, using language goals as semantic rewards to replace manual reward engineering in RL-based driving.
[xu2024vlm] 2024 VLM-AD LLaMA-2 NuScenes Vision, Text L2 Error, Collision Rate Reasoning
Innovation: Proposes VLM-AD, using VLMs as teachers to inject reasoning supervision into E2E driving without requiring them at inference.
[gopalkrishnan2024multi] 2024 EM-VLM4AD Multiple DriveLM Image, Text BLEU-4, METEOR Reasoning
Innovation: Presents a lightweight multi-frame VLM optimized for real-time autonomous driving with low memory and fast inference.
[sima2024drivelm] 2023 DriveLM-Agent BLIP-2 NuScenes Text, Image Accuracy Reasoning
Innovation: Introduces Graph VQA to enable multi-step, graph-structured reasoning in end-to-end driving via VLMs.
[li2023blip] 2023 BLIP-2 LLaMA-2-7B COCO, VQAv2 Text, Vision Accuracy, BLEU Captioning
Innovation: A two-stage strategy using frozen image and language models with a lightweight transformer for efficient vision-language pretraining.
[jia2024bev] 2024 BEV-CLIP CLIP+Llama2 NuScenes Text, BEV Accuracy, Captioning
Innovation: A retrieval method using LLMs and knowledge graphs for rich semantic scene matching and contextual reasoning capabilities.
[choudhary2024talk2bev] 2024 Talk2BEV BEVFormer+GPT-2 NuScenes Text, Image, LiDAR Accuracy, IoU Reasoning
Innovation: Introduces a LVLM interface for BEV maps enabling zero-shot visual and spatial reasoning without BEV-specific training.
[cheng2023language] 2023 MSSG GPT-2 NuScenes LiDAR, Text BEV AP, 3D AP Reasoning
Innovation: Proposes LiDAR Grounding with MSSG for efficient 3D language grounding in point clouds using joint LiDAR-language learning.
[wu2023referring] 2023 RMOT Swin-L+BERT KITTI Vision HOTA, DetA, Tracking
Innovation: A framework for tracking multiple objects in video using language-guided expressions and contextual scene-level understanding.
[chen2023clip2scene] 2023 CLIP2Scene CLIP+PointNet++ NuScenes, ScanNet Text, Vision mIoU, Accuracy Reasoning
Innovation: A framework to transfer CLIP knowledge to 3D scene understanding via semantic-driven contrastive learning.
[najibi2023unsupervised] 2023 CLIP ViT+Transformer Waymo Text, LiDAR 3D AP Detection, Reasoning
Innovation: Proposes an auto-labeling pipeline for open-set 3D perception using motion cues and vision-language distillation without human labels.
Wu et al. [wu2023referring] present a vision language task called Referring Multi-Object Referring (RMOT), which enables queries in natural language to refer to and track multiple objects over time in video sequences. RMOT supports flexible object cardinality and dynamic temporal variations in object status. Authors develop a benchmark, Refer-KITTI, based on the KITTI dataset, providing high-quality annotations with low labeling cost and supporting multi-object, temporally-aware queries. Additionally, their proposed TransRMOT uses an early-fusion cross-modal encoder and a query-decoupled decoder for simultaneous detection and tracking. TransRMOT performs better than CNN-based and Transformer-based baselines on the Refer-KITTI benchmark, highlighting the RMOT formulation’s effectiveness and the proposed architecture for real-world, language-guided multi-object tracking. The authors in [wu2025language] propose a large-scale benchmark to combine natural language prompts into 3D perception for autonomous driving. NuPrompt is based on the nuScenes dataset, which contains more than 40,000 finely annotated language-object pairs, with each prompt referring to multiple object trajectories across 3D, multiview, and temporal frames. For simplicity, the authors define a new task for predicting object motion based on descriptive prompts and propose a baseline model, PromptTrack, incorporating a novel prompt reasoning module into a transformer-based tracking architecture. The results show that PromptTrack outperforms several strong baselines and validates the effectiveness of prompt-guided object tracking and forecasting. Their method is the foundation for research at the intersection of vision-language modeling and AV systems.
Chen et al. [chen2023clip2scene] propose a framework, Contrastive Language Image Pretraining (CLIP) models, for 3D scene understanding tasks, particularly for semantic segmentation of point clouds. They address the challenges of limited 3D annotations and domain mismatch between 2D and 3D data. Also, the authors propose a Semantic-driven Cross-modal Contrastive Learning strategy combining semantic consistency and spatial-temporal consistency. This method enables effective annotation-free segmentation, achieving state-of-the-art results on datasets like nuScenes and ScanNet, and also shows better performance in few-shot learning scenarios. This work is among the first to bridge 2D vision language pre-training and 3D scene understanding through cross-modal contrastive learning. Taparia et al. [taparia2025vlc] present a fusion method that uses VLMs to improve object detection under various environmental conditions. Some fusion methods often apply static fusion rules and have challenges in unseen scenarios. At the same time, VLC Fusion addresses this limitation by conditioning the fusion process on environmental context extracted using large VLMs such as GPT-4o. The authors introduce an automated pipeline to extract and query application-specific ecological cues, which are then used to modulate multimodal feature fusion via Feature-wise Linear Modulation (FiLM). Experiments on Waymo Open (RGB + LiDAR) and ATR (visible + infrared) datasets show that VLC Fusion outperforms other baselines.

### v3

@@ 5.4 LLMs/VLMs-based Methods
Nowadays, there has been significant progress in LLMs/VLMs, and a wide variety of models have been developed for different domains and applications. Each model has different sizes, supported modalities, inference costs, and design architectures. Table 23 provides an in-depth overview of the most prominent LLM/VLM models, highlighting their developers, architectures, modalities, number of parameters, and key innovations.
Table 23: Summary and analysis of the significant LLM and VLM models, highlighting their advancements and key innovations.
Ref Year Name Developer Architecture Modalities Parameters Key Innovations
Large-Language Models (LLMs)
[52] 2025 Gemini 2.5 Google DeepMind Multimodal- Text, Vision $\sim$100 B $\square\checkmarkMultilingual$ Processing
Transformer $\square\checkmarkMultimodal$ Input
[13] 2025 Claude 4 Anthropic Constitutional- Text >>1 T $\square\checkmarkSafety$-focused LLM
Transformer $\square\checkmarkImproved$ Reasoning
[259] 2025 Grok 4 xAI GPT-style Text $\sim$1 B-100 B $\square\checkmarkTwitter$-integrated
Transformer $\square\checkmarkImproved$ Factuality
[100] 2025 Qwen-2.5 Alibaba Multimodal- Text, Vision +110 B $\square\checkmarkMultilingual$ Processing
Transformer $\square\checkmarkMultimodal$ Input
[173] 2024 GPT-o3 OpenAI GPT4-style Text, Vision >>1 T $\square\checkmarkMultimodal$ Input
Transformer $\square\checkmarkHigh$-speed
[64] 2024 LLaMA 3 Meta AI Decoder-only Text $\sim$8 B-70 B $\square\checkmarkImproved$ Scaling
Transformer $\square\checkmarkTraining$ Stability
[217] 2024 Gemini 1.5 Google DeepMind Multimodal- Text, Vision $\sim$100 B-1 T $\square\checkmarkContext$ Understanding
Transformer $\square\checkmarkMultilingual$ Processing
[9] 2024 Claude 3 Anthropic Constitutional- Text $\sim$100 B-1 T $\square\checkmarkVision$-language Integration
Transformer $\square\checkmarkEnhanced$ Reasoning
[153] 2024 DeepSeek-VL DeepSeek AI Vision-Language Text, Vision $\sim$1 B-100 B $\square\checkmarkMultimodal$ Tuning
Transformer $\square\checkmarkVision$-language Integration
[101] 2024 GPT-4o OpenAI Multimodal- Text, Audio $\sim$1.8 T $\square\checkmarkReal$-time Conversation
Transformer $\square\checkmarkEnd$-to-end Audio
[1] 2024 Phi-3 Microsoft Compact- Text $\sim$3.8 B $\square\checkmarkSmall$-scale
Transformer $\square\checkmarkCode$ Understanding
[221] 2023 LLaMA 2 Meta AI Decoder-only Text $\sim$7 B-70 M $\square\checkmarkChat$ Fine-tuning
Transformer $\square\checkmarkLogic$ Capabilities
[2] 2023 GPT-4 OpenAI Multimodal- Text, Vision >>1 T $\square\checkmarkMultimodal$ Reasoning
Transformer $\square\checkmarkVision$ Support
[8] 2023 PaLM 2 Google Parallel Layers Text $\sim$1 B-100 B $\square\checkmarkMultilingual$ Reasoning
Transformer $\square\checkmarkLogic$ Capabilities
[81] 2023 Falcon TII Auto-regressive Text $\sim$180 B $\square\checkmarkMassive$ scale MoE
Transformer $\square\checkmarkImproved$ Conversation
[188] 2019 GPT-2 OpenAI Decoder-only Text $\sim$1.5 B $\square\checkmarkGenerative$ Pretraining
Transformer $\square\checkmarkZero$-shot Generalization
[119] 2019 BART Meta AI Encoder–Decoder Text $\sim$406 M $\square\checkmarkBERT$-style Encoder
Transformer $\square\checkmarkGPT$-style Decoder
[58] 2018 BERT Google Bidirectional- Text $\sim$340 M $\square\checkmarkBidirectional$ Encoder
Transformer $\square\checkmarkMasked$ Modeling
Vision-Language Models (VLMs)
[160] 2025 CAKE Xiamen CLIP + Text, Vision $\sim$200 M $\square\checkmarkCategory$-aware Extraction
Faster R-CNN $\square\checkmarkNovel$-class Generalization
[316] 2025 LED Rutgers LLM-guided Text, Vision $\sim$230 M $\square\checkmarkFuses$ LLM Hidden States
Transformer $\square\checkmarkAvoids$ Synthetic Data
[49] 2024 YOLO-World Tencent Lab YOLO + Text, Vision $\sim$100 M $\square\checkmarkText$ Encoder Fusion
Text Encoder $\square\checkmarkTraining$-free Deployment
[144] 2024 Grounding-DINO IDEA DETR-style Text, Vision $\sim$172 M $\square\checkmarkLLM$-guided Region
Transformer $\square\checkmarkText$-guided Transformer
[254] 2023 CORA CPII CLIP Backbone + Text, Vision $\sim$200 M $\square\checkmarkCLIP$-based Anchor
DETR Transformer $\square\checkmarkText$-aware Proposal
[121] 2022 GLIP Microsoft Vision-Language Text, Vision $\sim$180 M $\square\checkmarkGrounded$ pretraining
Transformer $\square\checkmarkUnified$ VLM Backbone
[117] 2022 F-VLM Google CLIP-based T Text, Vision $\sim$400 M $\square\checkmarkFrozen$ CLIP Backbone
Transformer $\square\checkmarkLightweight$ Fine-tuning
[110] 2021 MDETR Facebook DETR-style Text, Vision $\sim$132 M $\square\checkmarkFirst$ Multimodal Detection
Transformer $\square\checkmarkLanguage$-modulated Queries
From Table 23, it is clear that LLM development has progressed significantly, from early models like GPT-1 and BERT to more advanced architectures such as GPT-4o, Gemini 2.5, and Claude 4. These models differ in architecture type, number of parameters, modality support, and their ability to perform reasoning, generation, or multimodal tasks. Some models, like GPT-3.5, GPT-4, and Claude, support tool-use and real-time applications, while others focus on domain-specific tasks. In the context of autonomous driving, LLMs can be integrated to enhance perception, facilitate natural language interaction, interpret sensor data, and support decision-making processes. Their ability to understand and generate human-like language enables more explainable and context-aware autonomous systems. These abilities make them powerful tools to complement traditional CV methods, especially when domain knowledge and semantic reasoning are needed. Meanwhile, the VLMs developed from early models like CLIP and MDETR to the latest advanced architectures such as Grounding-DINO and CAKE. While earlier models focused on contrastive learning and dual encoders, recent VLMs use advanced technologies such as LLM-guided prompts, grounding mechanisms, and hierarchical reasoning. These enhancements allow VLMs to better align visual and textual modalities, operate effectively in zero-shot settings, and support a wide range of vision-language tasks with higher accuracy and flexibility.